Gem Industries is a division of a major corporation. Last year the division had total sales of $24,048,000, net operating income of $2,765,520, and average operating assets of $6,012,000. The company's minimum required rate of return is 17%.

Required:

a. What is the division's margin? (Round your percentage answer to 2 decimal places.)

b. What is the division's turnover? (Round your answer to 2 decimal places.)

c. What is the division's return on investment (ROI)?
Answer #1

Margin=Net operating income/Sales

=2,765,520/24,048,000=11.5%

Turnover=Sales/Average operating assets

=24,048,000/6,012,000=4

ROI=Net operating income/Average operating assets

=2,765,520/6,012,000=46%